# Residual Diffusion Models for Variable-Rate Joint Source Channel Coding of MIMO CSI

### v1

@@ V-G Computational Complexity
We now evaluate the computational complexity of RD-JSCC by measuring the throughput of each module on both the COST2100 outdoor and 3GPP indoor datasets. Throughput is measured in terms of samples processed per second (samples/s), averaged over multiple runs with a batch size of $10^{3}$. All simulations were conducted on a system equipped with an AMD Ryzen Threadripper PRO 5975WX 32-Core processor and an NVIDIA GeForce RTX 4090 GPU.
@@
As shown in Tables III and IV, the encoder and decoder exhibit high throughput due to their lightweight convolutional architecture. The encoder consistently achieves throughput near $9.5\times 10^{4}$ samples/s across both datasets, making it suitable for low-power UEs. The decoder is slightly more complex due to residual layers, but still maintains a high throughput of approximately $8.25\times 10^{4}$ samples/s.
@@
In contrast, the diffusion model, though delivering significant performance improvements under complex channel conditions, introduces a higher computational cost. The 2-step residual diffusion refinement achieves throughput of $3.9\times 10^{3}$ samples/s on COST2100 and $8.2\times 10^{3}$ samples/s on the 3GPP indoor dataset. For applications requiring higher fidelity, the full 20-step diffusion yields throughput in the $10^{2}$ samples/s range, although the improvements in NMSE are marginal compared to 2-step diffusion. Hence, invoking the diffusion refinement module at the decoder can incur a throughput penalty of 10×10$\times$10 × to 100×100$\times$100 ×, and should therefore be used judiciously, only when necessary, based on the underlying channel complexity.
@@
These results reinforce the practicality of RD-JSCC’s hybrid decoding strategy, unlike a single diffusion-based solution presented in [35]. For simple channels, the decoder alone may suffice, while complex scenarios can selectively invoke diffusion-based refinement. Furthermore, early-exit and low-step inference modes offer flexible complexity-performance trade-offs.
Module Throughput (samples/s)
Encoder $9.5\times 10^{4}$
Decoder $8.2\times 10^{4}$
Diffusion (2-step) $3.9\times 10^{3}$
Diffusion (20-step) $3.9\times 10^{2}$
TABLE III: Throughput of modules in RD-JSCC for COST2100.
Module Throughput (samples/s)
Encoder $9.5\times 10^{4}$
Decoder $8.2\times 10^{4}$
Diffusion (2-step) $8.5\times 10^{3}$
Diffusion (20-step) $8.5\times 10^{2}$
TABLE IV: Throughput of modules in RD-JSCC for 3GPP indoor.
VI Variable Rate Compression
In practical wireless communication systems, the available uplink bandwidth for CSI feedback varies dynamically due to factors such as performance requirements and network congestion. Traditionally, deep learning-based compression methods address multiple compression rates by training separate models for each rate. However, this approach imposes significant overhead on resource-constrained user equipment (UE). To mitigate this, recent research has focused on developing single-model solutions capable of handling multiple compression rates.
For instance, CSI-PPPNet [52] introduces a one-sided architecture that supports arbitrary compression ratios using a single deep learning model. This design significantly reduces the number of parameters required at the UE, enhancing efficiency and simplifying deployment. Similarly, the Variable-Rate Code (VRC) approach dynamically allocates feedback bits based on estimated distortion, optimizing the trade-off between compression rate and reconstruction quality [26].
VI-A Matryoshka Representation Learning
Matryoshka Representation Learning (MRL), introduced in [53], is a representation learning framework designed to support multiple compression rates within a single model. The key idea behind MRL is to learn a single hierarchical representation that can be truncated at different lengths to yield compressed representations of varying fidelity. This is achieved by partitioning the learned representation vector into nested subsets, such that each prefix subset corresponds to a progressively higher compression rate. Importantly, MRL enforces joint optimization of all nested representations during training, enabling the model to perform well across a broad spectrum of rate-distortion trade-offs without the need for retraining or maintaining multiple models. This design allows for efficient and adaptive deployment in scenarios with dynamic bandwidth or storage constraints, while maintaining representational consistency and strong generalization across different rates.
$\mathcal{L}_{\text{MRL}}=\sum_{i=1}^{m}\lambda_{i}\|H_{d}-\hat{H}_{d}^{(k_{i})}\|^{2},$ (12)
where $\hat{H}_{d}^{(k_{i})}$ denotes the reconstructed CSI from the compressed representation of dimension $k_{i}$, and $\lambda_{i}$ are weighting factors that balance the importance across different rates. We choose $\lambda_{i}=1$ by default, assigning equal importance to all rates.
In Fig. 6, we show the results for the model optimized for 3 different feedback bandwidths $k=\{8,16,32\}$. We begin by choosing an encoder architecture to compress to a complex-valued latent dimension of $32$ and mask the appropriate elements using the MRL objective to create latents of dimensions of size $16$ and $8$. Based on available bandwidth, the encoder automatically chooses the best possible latent dimension, achieving an efficient rate-distortion trade-off across a range of feedback rates. We note that at a feedback bandwidth of $k=16$, we observe superior performance in Fig. 6 compared to Fig. 4. This improvement comes from the enhanced encoder complexity in terms of larger Dense layer required to accommodate the higher feedback bandwidth of $k=32$, which consequently benefits performance across all bandwidth configurations.
Figure 6: Optimizing for multiple feedback bandwidths using a single model with MRL objective. RD-JSCC dynamically adapts the latent dimensionality based on the available uplink bandwidth, achieving an efficient rate-distortion trade-off across a range of feedback rates.
@@ VII Practical considerations
While it is clear that diffusion models have the potential to enhance CSI reconstruction quality at the base station significantly, they also introduce substantial computational complexity. Therefore, it is crucial to justify this added complexity by employing diffusion-based refinement only when truly necessary. In the current state-of-the-art diffusion-based CSI compression method [35], the same network architecture is applied across both the Clustered Delay Line (CDL) and the more challenging COST2100 outdoor datasets, despite the latter being considerably more complex. Moreover, the comparisons to existing baselines do not account for the substantial differences in model size. For example, baseline models used in [35] such as CSINet and CRNet typically have around 400K parameters, whereas the generative diffusion-based approach presented utilizes approximately 15M parameters. This significant disparity in parameter count complicates the assessment of whether performance gains arise from the diffusion modeling itself or simply from increased model size.
@@
Architecture Choices.  We evaluate three different architectures for this study. First, we consider a simple convolutional autoencoder trained in a supervised manner. Second, we consider an autoencoder followed by a U-Net denoising network, also trained end-to-end with supervised loss. Finally, we evaluate the RD-JSCC model, where the autoencoder and U-Net are trained jointly using the residual diffusion objective. In all three cases, the architectures are appropriately scaled to maintain comparable parameter counts across models. Complete details of parameters used in each experiment are provided below, in Tab. V and Tab. VI.
Channel Model Choices.  In addition to the COST2100 outdoor scenario [50] analyzed in Sec. V, we now evaluate performance under the indoor open-area channel model specified in 3GPP TR 38.901 [51]. This environment is characterized by significantly lower spatial complexity and information density than COST2100, providing a useful contrast in dataset complexity. As seen in Sec. V, we first convert the spatial-frequency data into cropped angular-delay domain, preserving the first $32$ subcarriers. But, after decompression, we convert the estimated channel matrices from the cropped angular-delay domain back to the spatial-frequency domain and subsequently present the final NMSE results in the original SF domain.
@@
Notably, the flexible formulation of our residual diffusion framework provides an additional operational advantage, allowing for early termination of the decoding process after stage 1, regardless of prevailing channel conditions, thereby offering adaptive decoding complexity scaling depending on application requirements.
Model Encoder Decoder Denoising Model
DJSCC-CSINet+ (Small) 152K 118K
DJSCC-CSINet+ (Large) 152K 12M
ADJSCC (Small) 152K 118K
ADJSCC (Large) 152K 12M
U-Net-based JSCC 152K 118K 13.7M
RD-JSCC 152K 118K 13.9M
TABLE V: Parameter count for each model architecture used for COST2100 dataset experiment.
Model Encoder Decoder Denoising Model
DJSCC-CSINet+ (Small) 167K 191K
DJSCC-CSINet+ (Large) 167K 1.3M
ADJSCC (Small) 168K 197K
ADJSCC (Large) 168K 1.3M
U-Net-based JSCC 152K 118K 1.25M
RD-JSCC 152K 118K 1.28M
TABLE VI: Parameter count for each model architecture used for 3GPP indoor dataset experiment.
@@ VII-B Imperfect Channel Estimation
In the CSI compression literature, it is common to assume perfect CSI at UE. While this assumption is convenient for algorithm development and benchmarking, it rarely holds true in practical wireless communication systems. In realistic deployments, CSI must be acquired through pilot-based estimation at the receiver, which introduces errors due to noise, interference, and limited pilot resources. These estimation inaccuracies can significantly degrade the performance of both traditional and learning-based CSI compression schemes.
@@
In Fig. 8, we observe that RD-JSCC trained under the perfect CSI assumption suffers a substantial performance drop when exposed to severe channel estimation errors, particularly at a CNR of 00 dB. However, by incorporating channel estimation error directly into the training pipeline, RD-JSCC learns to operate robustly under imperfect CSI conditions, recovering most of the performance degradation.
Figure 8: By injecting channel estimation errors into the training pipeline, RD-JSCC becomes robust to imperfect CSI and recovers much of the performance degradation caused by estimation noise when tested at a CNR of 00 dB.
@@ VII-C Effect of Quantization
When using autoencoders for data compression, the encoder typically produces a continuous latent vector $\mathbf{s}\in\mathbb{C}^{k}$, which is often stored using 32-bit floating-point precision. This imposes a significant memory and communication overhead, especially in resource-constrained scenarios such as edge devices or embedded wireless systems. To reduce this burden, we utilize a companding-based quantization scheme that enables fixed-bit representation of latent vectors with minimal degradation in performance. Note that we quantize the real and imaginary values independently, using identical techniques.
@@
At inference, the quantized codes are dequantized and passed through the inverse $\mu$-law transformation:
@@
Figure 9: Effect of fixed-bit quantization on reconstruction quality. The encoder output can be quantized to as low as $4$ bits per dimension with negligible degradation in reconstruction performance.
@@ VIII-A Contribution of Residual Diffusion Formulation
To quantify the impact of the residual formulation in our denoising diffusion process, we conduct an ablation study comparing our proposed RD-JSCC scheme against a conventional GD-JSCC baseline. The baseline follows the standard formulation introduced in [35], where the reverse diffusion process is initialized from pure Gaussian noise and trained to generate the CSI purely based on the conditioning signal. In contrast, RD-JSCC initializes the reverse diffusion with a coarse reconstruction obtained from a lightweight autoencoder and performs iterative denoising on the residual between the ground-truth CSI and this initial estimate.
@@ VIII-B Contribution of Encoder
To investigate the role of encoder complexity in our architecture, we conducted ablation studies varying the encoder depth and channel capacity. Despite comprising a relatively small fraction of the overall model parameters, Fig. 11 demonstrates that encoder design significantly impacts reconstruction quality. When reducing our standard 4-layer encoder to a lighter 2-layer variant with fewer channels, performance deteriorates substantially—up to 2 dB loss in NMSE at equivalent compression rates. This confirms that the encoder, while computationally efficient, serves a critical function in extracting and preserving essential features from the input signal. These preserved features subsequently enable both the decoder and diffusion modules to achieve high-quality reconstruction.
Figure 11: Impact of encoder complexity on reconstruction performance. Reducing from a 4-layer to a 2-layer architecture with fewer channels degrades NMSE by up to 2 dB on the COST2100 dataset at a feedback bandwidth of $k=16$, despite the encoder’s minimal contribution to overall model complexity.
The exact architecture choices for the 2-layer and 4-layer encoder are provided in Tab. VII. The 2-layer encoder has a total of $66$K parameters, and the 4-layer encoder has a total of $152$K parameters. For reference, we recall from Tab. V that the full RD-JSCC model has $\approx 14$M parameters. Our findings suggest that appropriate encoder complexity represents an important design consideration that disproportionately influences overall system performance relative to its parameter count.
@@
Output Conv 2 (5,5)
TABLE VII: Comparison between 4-layer encoder and lightweight 2-layer encoder architectures used in ablation studies. The 4-layer configuration provides superior feature extraction capabilities, leading to better reconstruction performance as demonstrated in Fig. 11.
@@ IX Conclusion and Remarks
In this work, we introduce a hybrid JSCC scheme for MIMO CSI compression combining autoencoder-based initial estimation with diffusion-based refinement. Our residual diffusion approach initializes reverse diffusion with a coarse autoencoder estimate, tailoring the reverse diffusion specifically for reconstruction. This two-stage framework can be exited after either stage based on channel conditions and performance requirements, balancing computational complexity against fidelity. Further, using $\chi$-prediction, we enable single-step diffusion inference with minimal performance loss, substantially reducing latency. Our experiments across varying channel complexities show diffusion-based refinement delivers optimal value for complex channel distributions, while autoencoder solutions efficiently serve simpler channel distributions. These findings highlight diffusion models’ potential for enhancing CSI reconstruction at base stations under challenging uplink conditions. Future work includes developing multi-rate compression within a single model, investigating quantization effects, exploring weight quantization techniques, and optimizing the U-Net backbone for improved computational efficiency while preserving performance. Further, including additional loss terms such as the binary cross entropy loss between transmitted and received data or the effective MIMO channel capacity, to directly optimize for the downstream task [19] is also an interesting direction.

### v2

@@ VI-G Computational Complexity
We now evaluate the computational complexity of RD-JSCC by measuring the throughput of each module on both the COST2100 outdoor and 3GPP indoor datasets. Throughput is measured as samples processed per second (samples/s), averaged across multiple runs with a batch size of $10^{3}$. All simulations were conducted on a system equipped with an AMD Ryzen Threadripper PRO 5975WX 32-Core processor and a single NVIDIA GeForce RTX 4090 GPU. Training of the RD-JSCC model for the COST2100 dataset takes $\approx 15$ GPU hours, and the smaller RD-JSCC model trained on the 3GPP indoor dataset takes $\approx 6.8$ GPU hours.
@@
As shown in Table III, the encoder and decoder exhibit high inference throughput due to their lightweight convolutional architecture. The encoder consistently achieves throughput near $9.5\times 10^{4}$ samples/s across both datasets, making it suitable for low-power UEs. The decoder is slightly more complex due to residual layers, but still maintains a high throughput of approximately $8.25\times 10^{4}$ samples/s.
@@
In contrast, the diffusion model, though delivering significant performance improvements under complex channel conditions, introduces a higher computational cost. The 2-step residual diffusion refinement achieves throughput of $3.9\times 10^{3}$ samples/s on COST2100 and $8.2\times 10^{3}$ samples/s on the 3GPP indoor dataset. For applications requiring higher fidelity, the full 20-step diffusion yields throughput in the $10^{2}$ samples/s range, although the improvements in NMSE are marginal compared to 2-step diffusion. Hence, invoking the diffusion refinement module at the decoder can incur a throughput penalty of $10\times$ to $100\times$, and should therefore be used judiciously, only when necessary, based on the underlying channel complexity. Table III also shows the inference throughput using CPU, which shows $25-30\times$ degradation in performance.
@@
These results reinforce the practicality of RD-JSCC’s hybrid decoding strategy, unlike a single diffusion-based solution presented in [25]. For simple channels, the decoder alone may suffice, whereas in complex scenarios, diffusion-based refinement can be selectively invoked. Furthermore, early-exit and low-step inference modes offer flexible trade-offs between complexity and performance.
Module COST2100 3GPP Indoor
GPU CPU GPU CPU
Encoder $9.5\times 10^{4}$ $3.8\times 10^{3}$ $9.5\times 10^{4}$ $3.8\times 10^{3}$
Decoder $8.2\times 10^{4}$ $3.4\times 10^{3}$ $8.2\times 10^{4}$ $3.4\times 10^{3}$
Diffusion (2-step) $3.9\times 10^{3}$ $1.3\times 10^{2}$ $8.5\times 10^{3}$ $2.8\times 10^{2}$
Diffusion (20-step) $3.9\times 10^{2}$ $1.3\times 10^{1}$ $8.5\times 10^{2}$ $2.8\times 10^{1}$
TABLE III: GPU and CPU throughput (samples/s) of RD-JSCC modules for COST2100 and 3GPP indoor datasets. A lower-complexity channel requires a smaller model, thereby increasing throughput.
@@ VII Practical considerations
While diffusion models have the potential to significantly enhance CSI reconstruction quality at the base station, they also introduce substantial computational complexity. Therefore, it is crucial to justify this added complexity by employing diffusion-based refinement only when truly necessary. In the current state-of-the-art diffusion-based CSI compression method [25], the same network architecture is applied across both the Clustered Delay Line (CDL) and the more challenging COST2100 outdoor datasets, despite the latter being considerably more complex. Moreover, the comparisons to existing baselines do not account for the substantial differences in model size. For example, baseline models used in [25], such as CSINet and CRNet, typically have approximately 400K parameters, whereas the generative diffusion-based approach presented here utilizes approximately 15M parameters. This significant disparity in parameter count complicates assessing whether performance gains arise from the diffusion modeling itself or simply from increased model size.
@@
Architecture Choices. We evaluate three different architectures for this study. First, we consider a simple convolutional autoencoder trained in a supervised manner. Second, we consider an autoencoder followed by a U-Net denoising network, which is also trained end-to-end with a supervised loss. Finally, we evaluate the RD-JSCC model, where the autoencoder and U-Net are trained jointly using the residual diffusion objective. In all three cases, the architectures are appropriately scaled to maintain comparable parameter counts across models. Complete details of parameters used in each experiment are provided in Tab. IV and Tab. V.
Channel Model Choices. In addition to the COST2100 outdoor scenario [35] analyzed in Sec. VI, we now evaluate performance under the indoor open-area channel model specified in 3GPP TR 38.901 [49]. This environment is characterized by significantly lower spatial complexity and information density than COST2100, providing a useful contrast in dataset complexity. As seen in Sec. VI, we first convert the spatial-frequency data into the cropped angular-delay domain, preserving the first $32$ subcarriers. But, after decompression, we convert the estimated channel matrices from the cropped angular-delay domain back to the spatial-frequency domain and subsequently present the final NMSE results in the original SF domain.
@@
Notably, the flexible formulation of our residual diffusion framework provides an additional operational advantage: it allows early termination of the decoding process after stage 1, regardless of prevailing channel conditions, thereby enabling adaptive decoding-complexity scaling based on application requirements.
[…]
DJSCC-CSINet+ (Small) 152K 21.3M 118K 26.4M
DJSCC-CSINet+ (Large) 152K 21.3M 12M 320M
ADJSCC (Small) 152K 21.3M 118K 26.4M
ADJSCC (Large) 152K 21.3M 12M 320M
U-Net-based JSCC 152K 21.3M 118K 26.4M 13.7M 135.6M
RD-JSCC (2-step) 152K 21.3M 118K 26.4M 13.9M 271.8M
TABLE IV: Parameter counts and FLOPs comparison across methods for COST2100 dataset.
[…]
DJSCC-CSINet+ (Small) 167K 27.8M 191K 35.7M
DJSCC-CSINet+ (Large) 167K 27.8M 1.3M 63.4M
ADJSCC (Small) 168K 28.1M 197K 35.9M
ADJSCC (Large) 168K 28.1M 1.3M 63.7M
U-Net-based JSCC 152K 21.3M 118K 26.4M 1.25M 29.5M
RD-JSCC (2-step) 152K 21.3M 118K 26.4M 1.28M 59.7M
TABLE V: Parameter counts and FLOPs comparison across methods for 3GPP Indoor dataset.
@@ VII-B Imperfect Channel Estimation
In the CSI compression literature, it is common to assume perfect CSI at UE. While this assumption is convenient for algorithm development and benchmarking, it rarely holds true in practical wireless communication systems. In realistic deployments, CSI must be acquired via pilot-based estimation at the receiver, which introduces errors arising from noise, interference, and limited pilot resources. These estimation inaccuracies can significantly degrade the performance of both traditional and learning-based CSI compression schemes.
@@
In Fig. 9, we observe that RD-JSCC trained using error-free CSI under the perfect CSI assumption suffers a substantial performance drop when exposed to severe channel estimation errors, particularly at a CNR of 0 dB. However, by incorporating channel estimation error directly into the training pipeline, RD-JSCC learns to operate robustly under imperfect CSI, recovering most of the performance loss.
Figure 9: By injecting channel estimation errors into the training pipeline, RD-JSCC becomes robust to imperfect CSI and recovers much of the performance degradation caused by estimation noise when tested at a CNR of 0 dB.
@@ VII-C Effect of Quantization
When using autoencoders for data compression, the encoder typically produces a continuous latent vector $\mathbf{s}\in\mathbb{C}^{k}$, which is often stored using 32-bit floating-point precision. This imposes significant memory and communication overhead, particularly in resource-constrained settings, such as edge devices and embedded wireless systems. To reduce this burden, we use a companding-based quantization scheme that enables fixed-bit representation of latent vectors with minimal performance degradation. Note that we quantize the real and imaginary values independently, using identical techniques.
@@
At inference, the quantized codes are de-quantized and passed through the inverse $\mu$-law transformation:
@@
Figure 10: Effect of fixed-bit quantization on reconstruction quality. The encoder output can be quantized to as few as 4 bits per dimension with negligible degradation in reconstruction performance.
@@ VIII-A Contribution of Residual Diffusion Formulation
To quantify the impact of the residual formulation in our denoising diffusion process, we conduct an ablation study comparing our proposed RD-JSCC scheme against a conventional Gaussian Diffusion (GD) JSCC baseline. The baseline follows the standard formulation introduced in [25], where the reverse diffusion process is initialized from pure Gaussian noise and trained to generate the CSI purely based on the conditioning signal. In contrast, RD-JSCC initiates reverse diffusion with a coarse reconstruction from a lightweight autoencoder and iteratively denoises the residual between the ground-truth CSI and this initial estimate.
@@ VIII-B Contribution of Encoder
To investigate the role of encoder complexity in our architecture, we conducted ablation studies varying the encoder depth and channel capacity. Despite comprising a relatively small fraction of the overall model parameters, Fig. 12 demonstrates that encoder design significantly impacts reconstruction quality. When reducing our standard 4-layer encoder to a lighter 2-layer variant with fewer channels, performance deteriorates substantially, by up to 1 dB loss in NMSE at equivalent compression rates. This confirms that the encoder, while computationally efficient, serves a critical function in extracting and preserving essential features from the input signal. These preserved features subsequently enable both the decoder and diffusion modules to achieve high-quality reconstruction.
Figure 12: Impact of encoder complexity on reconstruction performance. Reducing from a 4-layer to a 2-layer architecture with fewer channels degrades NMSE by up to 1 dB on the COST2100 dataset at a feedback bandwidth of $k=16$, despite the encoder’s minimal contribution to overall model complexity.
The exact architecture choices for the 2-layer and 4-layer encoder are provided in Tab. VI. The 2-layer encoder has a total of 6666K parameters, and the 4-layer encoder has a total of 152152K parameters. For reference, we recall from Tab. IV that the full RD-JSCC model has $\approx 14$M parameters. Our findings suggest that appropriate encoder complexity represents an important design consideration that disproportionately influences overall system performance relative to its parameter count.
@@
Output Conv 2 (5,5) 2 (7,7)
TABLE VI: Comparison between 4-layer encoder and lightweight 2-layer encoder architectures used in ablation studies. The 4-layer configuration provides superior feature extraction capabilities, leading to better reconstruction performance as demonstrated in Fig. 12.
VIII-C Contribution of SNR Adaptation
We investigate the contribution of SNR side information by training RD-JSCC without the SNR adaptation module and comparing performance across uplink SNR conditions. We see from Fig. 12 that SNR conditioning provides consistent improvements across all SNR values, with up to $0.35$ dB NMSE gain in the high SNR regime and up to $0.1$ dB gain in the lower-SNR region, demonstrating that SNR-aware denoising enables the diffusion model to adapt its reconstruction strategy based on channel quality rather than learning a single solution.
Figure 13: Impact of SNR side-information on RD-JSCC performance. Removing the SNR adaptation module results in an NMSE degradation of up to $0.35$ dB for a given SNR.
@@ IX Conclusion and Remarks
In this work, we introduce a hybrid JSCC scheme for MIMO CSI compression that combines an autoencoder-based initial estimation with diffusion-based refinement. Our residual diffusion approach initializes reverse diffusion with a coarse autoencoder estimate, thereby tailoring it specifically for reconstruction. This two-stage framework can be exited after either stage based on channel conditions and performance requirements, balancing computational complexity against fidelity. Further, using $\chi$-prediction, we enable a low-complexity 2-step diffusion inference with minimal performance loss, substantially reducing latency. Our experiments across varying channel complexities show diffusion-based refinement delivers optimal value for complex channel distributions, while autoencoder solutions efficiently serve simpler channel distributions. These findings highlight the potential of diffusion models to enhance CSI reconstruction at base stations under challenging uplink conditions. Future work includes developing multi-rate compression within a single model, investigating quantization effects, exploring weight-quantization techniques, and optimizing the U-Net backbone to improve computational efficiency while preserving performance. Further, including additional loss terms, such as the binary cross-entropy loss between transmitted and received data or the effective MIMO channel capacity, to directly optimize for the downstream task [31] is also an interesting direction.
Acknowledgment
The authors thank Charlie Zhang for insightful discussions. This work was partly supported by Samsung Research America through the 6G@@UT center within the Wireless Networking and Communications Group (WNCG) at the University of Texas at Austin, ARO Award W911NF2310062, ONR Award N000142412542, and NSF Award 2443857.